# Fluid Antenna-enabled Near-Field Integrated Sensing, Computing and Semantic Communication for Emerging Applications

### v1

@@ IV-C FA Position Optimisation
The optimisation problem in (33) is non-convex due to the non-convex nature of the objective function in (33a). To address this, we first describe a benchmark algorithm widely used in the literature, such as in [54, 55, 56]. We then propose a new algorithm that reduces computational complexity while maintaining performance.
@@ IV-C1 Benchmark method (second-order Taylor expansion)
Likewise, the upper bound of $R_{l|k}$ is determined by finding the lower bound of $\log_{2}\left(C_{l|k}\right)$ and the upper bound of $\log_{2}\left(D_{l|k}\right)$, as shown in (37) and (38), respectively. In this manner, the constraint in (33b) is replaced by the objective in (39), with the subscript $e$ indicating the number of epochs. To ensure consistency in matrix dimensions, the Hessian matrices are replaced by scalar values $\epsilon$ and $\delta$. The selection of $\epsilon$ and $\delta$ in (35)–(38) should be made to avoid the appearance of $\mathbf{u}\leq\mathbf{u}_{e}$, which would result in a negative gradient that minimises the objective function instead of maximising it. A recommended approach is to ensure that the conditions $\epsilon_{k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(A_{k,e}\right)$, $\delta_{k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(B_{k,e}\right)$, $\epsilon_{l|k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(C_{l|k,e}\right)$, and $\delta_{l|k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(D_{l|k,e}\right)$ hold, as shown in [57, 58]. One practical choice for $\epsilon$ and $\delta$ is to use the maximum eigenvalues of the respective matrices, ensuring that the conditions are satisfied while also improving the stability of the optimisation process.
@@
$\log_{2}\left(A_{k}\right)\geq\log_{2}\left(A_{k,e}\right)+\nabla\log_{2}\left(A_{k,e}\right)\left(\mathbf{u}-\mathbf{u}_{e}\right)^{T}-\left(\mathbf{u}-\mathbf{u}_{e}\right)\frac{\nabla^{2}\log_{2}\left(A_{k,e}\right)}{2}\left(\mathbf{u}-\mathbf{u}_{e}\right)^{T},\quad$ (35)
@@
By substituting each Hessian matrix in (39) with the corresponding scalar values $[\delta_{k},\epsilon_{k},\delta_{l|k},\epsilon_{l|k}]$, the optimisation problem in (33) becomes convex and can be efficiently solved using existing optimisation tools. The computational complexity at each epoch is $\mathcal{O}\left(\left(2N_{\text{tx}}N_{\text{tz}}\right)^{3}\right)$.
@@ IV-C2 Projected BFGS
Lemma 1 (BFGS [59]).
@@
×$\;\times\bigg{(}\mathbf{I}-\frac{\Delta\mathbf{u}_{e+1}^{T}\Delta\left(\nabla_{e+1}\right)}{\Delta\left(\nabla_{e+1}\right)\Delta\mathbf{u}_{e+1}^{T}}\bigg{)}+\frac{\Delta\mathbf{u}_{e+1}^{T}\Delta\mathbf{u}_{e+1}}{\Delta\left(\nabla_{e+1}\right)\Delta\mathbf{u}_{e+1}^{T}},$
@@ V Simulation Results
In this section, we present numerical results to evaluate the effectiveness of the proposed design. The default values for the parameters are summarised in Table I. The simulation utilises the mmWave frequency band, where the small antenna size leads to narrower movable regions for the FAs compared to existing works, such as [60, 55]. For simplicity, we assume that the movable region of the FAs is square-shaped. Consequently, the movable range along the x-axis or z-axis is $\sqrt{C_{i}}$. While the analysis can be extended to non-square movable regions, this falls beyond the scope of the current study.
@@ V-B Semantic Communication Performance
Fig. 7 illustrates the relationship between the movable range of FAs, denoted as $\sqrt{C_{i}}$, and the resulting per antenna gain compared with the case of using FPAs, expressed as a percentage. Three methods are compared: the benchmark method, the projected BFGS method, and randomised FAs, where the positions of FAs are randomly generated within the movable range. The result reveals that both the benchmark method and the projected BFGS benefit from increasing the movable range $\sqrt{C_{i}}$, as larger movable regions allow for more exploration of better antenna positioning, leading to significant performance gains. For example, when the movable range is 20 cm, the per antenna gain for the benchmark method is 0.53% while the projected BFGS gives 0.56%. Notably, the projected BFGS consistently outperforms the benchmark method across all $\sqrt{C_{i}}$ values, and the difference in the performance becomes more notable as $\sqrt{C_{i}}$ increases. This demonstrates the superior optimisation capabilities of the projected BFGS algorithm, on top of its lower complexity when compared with the benchmark method. In contrast, the random method exhibits a nearly flat and negligible performance gain. For instance, with 20cm of movable range, the per antenna gain is only 0.04%, which is ten times less than the other two algorithms can achieve. The figure also compares the projected BFGS algorithm with and without using semantic. Notably, incorporating semantic processing consistently enhances the per-antenna gain by approximately 1.2–1.5 times compared to the non-semantic design (i.e., conventional ISAC designs as in [14, 61]). This clearly highlights the advantage of integrating semantic into the model.
@@ V-C Sensing Performance
[…]
@@ Appendix A Proof and Derivation of Equation (15)
Utilising the established result for linear Gaussian models as detailed in [62, Chapter 4], the FIM is explicitly given by
@@ Appendix C Convergence Analysis
The sub-problem $M_{1}$ is non-convex with respect to $\mathbf{w}_{k}$ and $\mathbf{R}_{x}$. We show that $M_{1}$ increases when jointly updating $[\mathbf{w}_{k},\mathbf{R}_{x},f_{l}]$:
@@
The sub-problem $M_{2}$ is non-convex with respect to $\mathbf{u}$. The projected BFGS algorithm, combined with backtracking line search, generates a sequence of iterates $\{\mathbf{u}_{e}\}$ to maximise $g(\mathbf{u})$. Since $g(\mathbf{u})$ is twice Lipschitz continuously differentiable and locally quadratic in $\mathbf{u}$, and the approximated Hessian $(\nabla^{2}_{e})^{-1}$ remains positive definite, the algorithm is guaranteed to converge to a stationary point under standard conditions [63, 64].

### v2

@@ IV-C FA Position Optimisation
The optimisation problem in (33) is non-convex due to the non-convex nature of the objective function in (33a). To address this, we first describe a benchmark algorithm widely used in the literature, such as in [ma2024multi, feng2024weighted, lyu2024flexible]. We then propose a new algorithm that reduces computational complexity while maintaining performance.
@@ IV-C1 Benchmark method (second-order Taylor expansion)
Likewise, the upper bound of $R_{l|k}$ is determined by finding the lower bound of $\log_{2}\left(C_{l|k}\right)$ and the upper bound of $\log_{2}\left(D_{l|k}\right)$, as shown in (37) and (38), respectively. In this manner, the constraint in (33b) is replaced by the objective in (39), with the subscript $e$ indicating the number of epochs. To ensure consistency in matrix dimensions, the Hessian matrices are replaced by scalar values $\epsilon$ and $\delta$. The selection of $\epsilon$ and $\delta$ in (35)–(38) should be made to avoid the appearance of $\mathbf{u}\leq\mathbf{u}_{e}$, which would result in a negative gradient that minimises the objective function instead of maximising it. A recommended approach is to ensure that the conditions $\epsilon_{k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(A_{k,e}\right)$, $\delta_{k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(B_{k,e}\right)$, $\epsilon_{l|k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(C_{l|k,e}\right)$, and $\delta_{l|k}\mathbf{I}\succeq\nabla^{2}\log_{2}\left(D_{l|k,e}\right)$ hold, as shown in [sun2016majorization, boyd2004convex]. One practical choice for $\epsilon$ and $\delta$ is to use the maximum eigenvalues of the respective matrices, ensuring that the conditions are satisfied while also improving the stability of the optimisation process.
@@
$\log_{2}\left(A_{k}\right)\geq\log_{2}\left(A_{k,e}\right)+\nabla\log_{2}\left(A_{k,e}\right)\left(\mathbf{u}-\mathbf{u}_{e}\right)^{T}-\left(\mathbf{u}-\mathbf{u}_{e}\right)\frac{\nabla^{2}\log_{2}\left(A_{k,e}\right)}{2}\left(\mathbf{u}-\mathbf{u}_{e}\right)^{T},\qquad$ (35)
@@
By substituting each Hessian matrix in (39) with the corresponding scalar values $\delta_{k},\epsilon_{k},\delta_{l|k},\epsilon_{l|k}]$, the optimisation problem in (33) becomes convex and can be efficiently solved using existing optimisation tools. The computational complexity at each epoch is $\mathcal{O}\left(\left(2N_{\text{tx}}N_{\text{tz}}\right)^{3}\right)$.
@@ IV-C2 Projected BFGS
  Lemma 1 (BFGS [fletcher2000practical]).
@@
$\;\times\bigg(\mathbf{I}-\frac{\Delta\mathbf{u}_{e+1}^{T}\Delta\left(\nabla_{e+1}\right)}{\Delta\left(\nabla_{e+1}\right)\Delta\mathbf{u}_{e+1}^{T}}\bigg)+\frac{\Delta\mathbf{u}_{e+1}^{T}\Delta\mathbf{u}_{e+1}}{\Delta\left(\nabla_{e+1}\right)\Delta\mathbf{u}_{e+1}^{T}},$
@@ V Simulation Results
In this section, we present numerical results to evaluate the effectiveness of the proposed design. The default values for the parameters are summarised in Table I. The simulation utilises the mmWave frequency band, where the small antenna size leads to narrower movable regions for the FAs compared to existing works, such as [qin2024cramer, feng2024weighted]. For simplicity, we assume that the movable region of the FAs is square-shaped. Consequently, the movable range along the x-axis or z-axis is $\sqrt{C_{i}}$. While the analysis can be extended to non-square movable regions, this falls beyond the scope of the current study.
@@ V-B Semantic Communication Performance
Fig. 7 illustrates the relationship between the movable range of FAs, denoted as $\sqrt{C_{i}}$, and the resulting per antenna gain compared with the case of using FPAs, expressed as a percentage. Three methods are compared: the benchmark method, the projected BFGS method, and randomised FAs, where the positions of FAs are randomly generated within the movable range. The result reveals that both the benchmark method and the projected BFGS benefit from increasing the movable range $\sqrt{C_{i}}$, as larger movable regions allow for more exploration of better antenna positioning, leading to significant performance gains. For example, when the movable range is 20 cm, the per antenna gain for the benchmark method is 0.53% while the projected BFGS gives 0.56%. Notably, the projected BFGS consistently outperforms the benchmark method across all $\sqrt{C_{i}}$ values, and the difference in the performance becomes more notable as $\sqrt{C_{i}}$ increases. This demonstrates the superior optimisation capabilities of the projected BFGS algorithm, on top of its lower complexity when compared with the benchmark method. In contrast, the random method exhibits a nearly flat and negligible performance gain. For instance, with 20cm of movable range, the per antenna gain is only 0.04%, which is ten times less than the other two algorithms can achieve. The figure also compares the projected BFGS algorithm with and without using semantic. Notably, incorporating semantic processing consistently enhances the per-antenna gain by approximately 1.2–1.5 times compared to the non-semantic design (i.e., conventional ISAC designs as in [zhao2024modeling, hao2024fluid]). This clearly highlights the advantage of integrating semantic into the model.
@@ V-C Sensing Performance
Fig. 10 shows the achieved MSE compared to the CRB upper limit ($\xi$). The MSE is calculated using maximum likelihood estimation (MLE). The ideal reference line represents the case where $\text{MSE}=\text{CRB}$, corresponding to the tight theoretical lower bound for a linear model. As $\xi$ increases from 0.3 to 1.0, a noticeable gap appears between the actual MSE and the ideal MSE, indicating that while the estimator closely approaches the theoretical bound, practical limitations, such as the limited number of frames ($F=100$), introduce a deviation of approximately 10%, consistent with the standard error ($1/\sqrt{F}=10\%$). Despite this, the relatively small and stable gap across the entire range of $\xi$ underscores the robustness and effectiveness of the proposed design under varying estimation accuracy requirements.
@@ Appendix A Proof and Derivation of Equation (15)
Utilising the established result for linear Gaussian models as detailed in [kay1993fundamentals, Chapter 4], the FIM is explicitly given by
@@ Appendix C Convergence Analysis
The sub-problem $M_{1}$ is non-convex with respect to $\mathbf{w}_{k}$ and $\mathbf{R}_{x}$. We show that $M_{1}$ increases when jointly updating $\mathbf{w}_{k},\mathbf{R}_{x},f_{l}]$:
@@
The sub-problem $M_{2}$ is non-convex with respect to $\mathbf{u}$. The projected BFGS algorithm, combined with backtracking line search, generates a sequence of iterates $\{\mathbf{u}_{e}\}$ to maximise $g(\mathbf{u})$. Since $g(\mathbf{u})$ is twice Lipschitz continuously differentiable and locally quadratic in $\mathbf{u}$, and the approximated Hessian $(\nabla^{2}_{e})^{-1}$ remains positive definite, the algorithm is guaranteed to converge to a stationary point under standard conditions [schmidt2009optimizing, nocedal1999numerical].